40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 19 14:06:36 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_14-06-11.png
    Reply  Fri Apr 19 20:21:52 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_20-27-04.png
       Reply  Wed Apr 24 16:06:44 2019, gautam, Update, PSL, Innolight NPRO shutoff PMCrefl.png
          Reply  Thu Apr 25 01:03:29 2019, gautam, Update, PSL, MC trans is ~1000 cts (~7%) lower than usual 
          Reply  Thu Apr 25 17:31:56 2019, gautam, Update, PSL, Innolight NPRO shutoff EdwinShutoff20190425.pngEdwinShutdown_zoomOut.png
             Reply  Mon Apr 29 16:25:22 2019, gautam, Update, PSL, PSL turned on again 
                Reply  Fri May 3 12:51:58 2019, gautam, Update, PSL, PSL turned on again 
                   Reply  Mon May 6 10:51:43 2019, gautam, Update, PSL, PSL turned off again NPROshutoff.png
                   Reply  Wed May 8 19:04:20 2019, rana, Update, PSL, PSL turned on again 
                      Reply  Thu May 9 19:50:04 2019, gautam, Update, PSL, PSL turned off again Screenshot_from_2019-05-09_19-49-29.png
                         Reply  Mon May 13 10:45:38 2019, gautam, Update, PSL, PSL turned ON again IMG_7431.JPG
                            Reply  Wed May 15 10:56:47 2019, gautam, Update, PSL, PSL turned ON again 
Message ID: 14609     Entry time: Wed May 15 10:56:47 2019     In reply to: 14605
Author: gautam 
Type: Update 
Category: PSL 
Subject: PSL turned ON again 

To test the hypothesis that the fan replacement had any effect on the NPRO shutoff phenomena, I turned the HEPA on the PSL table down to the nominal 30% setting at ~10am.

Tomorrow I will revert the laser crystal temperature to whatever the nominal value was. If the NPRO runs in that configuration (i.e. the only change from March 2019 are the diode TEC setpoints and the new fan on the back of the controller), then hurray.

ELOG V3.1.3-